OpenStudy (anonymous):

Will Fan and Medal Write an equation of a circle with the given center and radius. center (-7, -6) radius 2 a (x+7)^2+(y+6)^2=4 b (x-7)^2+(y-6)^2=2 c (x+7)^2+(y+6)^2=2 d (x-7)^2+(y-6)^2=4

OpenStudy (anonymous):

are you good at math? @golden_bullets

OpenStudy (anonymous):

@iTiaax @jdoe0001

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/example-2-equation-of-circle.png <-- you are given the radius and center, thus just plug your values in

OpenStudy (anonymous):

@jdoe0001 explain

OpenStudy (jdoe0001):

\(\bf center \quad ({\color{red}{ -7}}, {\color{blue}{ -6}}) \qquad radius=2 \\ \quad \\ (x-{\color{red}{ h}})^2+(x-{\color{blue}{ k}})^2=r^2\)

OpenStudy (jdoe0001):

so... what would you get?

OpenStudy (anonymous):

okay, (x+7) ^2+(x+6)^6=4

OpenStudy (jdoe0001):

ooops... darn that should be a "y" so \(\bf center \quad ({\color{red}{ -7}}, {\color{blue}{ -6}}) \qquad radius=2 \\ \quad \\ (x-{\color{red}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2\)

OpenStudy (anonymous):

@jdoe0001 what now

OpenStudy (jdoe0001):

plug the values in

OpenStudy (anonymous):

didn't I do that? (x+7) ^2+(y+6)^6=4 @jdoe0001 how is this and now what?